I am working on a math homework sheet. I am having trouble with some problems. Can someone please help me figure these out.

Find the equation of the parabola by the given information.

13.) Vertex at the origin, focus at (0,3).

Answer by scott8148(5896)

a general equation for a parabola is (x-h)^2=4p(y-k)

(h,k) is the vertex, and p is the distance from the vertex to the focus

substituting __ (x-0)^2=4(3-0)(y-0) __ x^2=12y